Hello,
I have worked out a story problem but I do not know if I am on the right track with it. Here is the problem:
The period T ( time in seconds for one complete cycle) of a simple pendulum is related tot he lenght L ( in feet) of the pendulum by the formula 8T^2=pie^2L. If a child is on a swing with a 10-ft chain, then how long does it take to complete one cycle of the swing?
This is how I worked it out:
T~2pie(46)
L is length of pendulum in meters
G is grauitational acceralation=9.8m/s.
10'=3.05m
t=2pie(l/G)
T=2pie (3.05/9.80=3.5 seconds is my answer

Answer by stanbon(75887)

The period T ( time in seconds for one complete cycle) of a simple pendulum is related the lenght L ( in feet) of the pendulum by the formula 8T^2=pie^2L. If a child is on a swing with a 10-ft chain, then how long does it take to complete one cycle of the swing?
--------------------
8T^2=pie^2*L
8T^2 = pi^2 *10
t^2 = (10/8)pi^2
t^2 = (5/4)pi^2
t = [sqrt(5)/2]pi
t = 1.1180..* pi seconds
====================
Cheers,
Stan H.